OpenStudy (anonymous):

2/3 x N = 12 Find the value for N thank you for helping me

jimthompson5910 (jim_thompson5910):

2/3 x N = 12 2 x N = 12 x 3 2 x N = 36 N = 36/2 N = 18

OpenStudy (anonymous):

thank u so much.. that is the answer i got!!